OpenStudy (anonymous):

Hello, I need help sloving a pythagorean problem. The side of the triangle is 5m,3m, and 2m+6

OpenStudy (anonymous):

the formula is a^2+b^2=c^2

OpenStudy (anonymous):

Which one is the hypotnus? the longest side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

rigth i'm using that formula but my answers are coming up negative? :-(

OpenStudy (anonymous):

are you trying to solve for m?

OpenStudy (anonymous):

no romero

OpenStudy (anonymous):

yes slove for m

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got negative 5 and 1 and then i time them numbers by the giving number and they are still negative

OpenStudy (anonymous):

\[(5m) + (3m)^2 = (2m+ 6)^2\]

OpenStudy (anonymous):

woops forgot a squared

OpenStudy (anonymous):

(5m)^2 + (3m)^2 = (2m + 6)^2 m = -0.76619 or m = 1.56619 But m = -0.76619 is an extraneous solution, so m = 1.56619 Consequently, the sides are about 7.83, 4.70, and 9.13

OpenStudy (anonymous):

\[(5m)^2 + (3m)^2 = (2m+6)^2\]

OpenStudy (anonymous):

Romero correct....dinosaur that looks correct how you come up with that answer?

OpenStudy (anonymous):

You have a quadratic so you can have two values for m

OpenStudy (anonymous):

if square all the numbers you get\[25^2 + 9m^2= 4m^2 + 24m+ 36\] get everything to one side you have \[30m^2 - 24m -36 = 0\]

OpenStudy (anonymous):

where a = 30 b =-24 c=-36 and you use the quadratic formula to find the answers. Do you know to do from now on?

OpenStudy (anonymous):

Romero that is what i got too and at the end the number are not coming out as whole numbers

OpenStudy (anonymous):

Does the question say anything about what side is the longest? or m belongs to some interval?

OpenStudy (anonymous):

the picture shows 5m to be the longest side

OpenStudy (anonymous):

I see!

OpenStudy (anonymous):

lol wow then we have it wrong

OpenStudy (anonymous):

I thought so, because that would make the answer look nicer :)

OpenStudy (anonymous):

if 5m is the longest side, then: \[(5m)^2=(3m)^2+(2m+6)^2 \implies 25m^2=9m^2+4m^2+24m+36\] \[\implies 12m^2-24m-36=0 \implies m^2-2m-3=0\] then after factorization: \[(m-3)(m+1)=0 \implies m=3, m=-1\]

OpenStudy (anonymous):

since the value of m represents length, we only take the positive value, that's m=3.

OpenStudy (anonymous):

awww thanks that look like that might be answer your a lifesaver :-)

OpenStudy (anonymous):

then your sides are: 5m=15, 3m=9 and 2m+6=12

OpenStudy (anonymous):

You're welcome :)